# Are We There Yet? A Brief Survey of Music Emotion Prediction Datasets, Models and Outstanding Challenges

### v2

@@ IV Models and approaches
In this section, we briefly touch upon some of the more recent MER models, highlighting the state-of-the-art approaches over the last few years. This is not an exhaustive overview, and for a more comprehensive review, readers are referred to other survey papers [11, 12, 13, 9, 14]. The aim of this section is to point out the current state-of-the-art and various approaches over the last five years. The models were selected through Google Scholar searches using search terms ‘music emotion prediction model’, ‘affective music prediction model’, and ‘music emotion recognition model’ from dates 2020, as well as by following links in the articles.
@@
Figure 1: Number of Models that use Datasets in Music Emotion Recognition Models since 2020.
@@
In recent years, various deep-learning MER models have been developed, employing techniques such as Convolutional Neural Networks (CNNs) [77, 79, 81, 82, 83, 72, 91, 84, 85, 86, 97, 104, 50, 100, 107, 110, 109], Recursive Neural Networks (RNNs) such as Gated Recurrent Units (GRUs) [100] and Long-Short Term Memory networks (LSTMs) [77, 78, 79, 80, 84, 85, 87, 97, 46, 50, 106], and more recently, Transformer architectures [88, 105, 50, 102]. Additionally, generative models such as Generative Adversarial Networks (GANs) have been applied. For instance, Huang et al. [90] developed a GAN-based model for emotion recognition using audio inputs under IoT environments.
@@
On the Jamendo dataset, several models have demonstrated noteworthy performance. For example, Mayerl et al. [73] employed a CNN-based (ResNet18) model using clustering-based ensembles, achieving a PR-AUC-macro score of 0.1087 and a ROC-AUC-macro score of 0.7047. Tan et al. [74], using a semi-supervised learning method and a CRNN architecture, reported higher scores of PR-AUC-macro 0.1357 and ROC-AUC-macro 0.7687. Pham et al. [86] achieved similar results with an EfficientNet-based CNN model, utilizing a co-teaching strategy to manage noisy data, resulting in PR-AUC-macro 0.1435 and ROC-AUC-macro 0.7599. Bour et al. [75] leveraged frequency-dependent convolutions in a CNN model to achieve the best performance with PR-AUC-macro 0.1509 and ROC-AUC-macro 0.7748, the highest score at MediaEval 2021.
@@
Notably, hybrid architectures combining multiple deep learning techniques, such as CNN-LSTM models, have shown promising results. For instance, Chen and Li [79] employed CNN-LSTM models for emotion recognition, outperforming standalone CNN and LSTM models. They achieved the highest accuracy of 68.1% for audio classification and 74.2% for lyric classification on the Last.fm dataset. Transformer-based models have also recently gained traction, such as those by Suresh et al. [105] and Agrawal et al. [88], which use Transformer architectures to improve MER performance. For example, the Transformer-based model proposed by Suresh et al. [105] outperformed other models, such as CNN and Bi-GRU, achieving the highest accuracy of 77.94% on the MoodyLyrics dataset.
@@
Other models have employed more traditional machine learning techniques alongside deep learning approaches. Medina et al. [98] focused on classifying emotions from audio using SVM, Random Forest, and Multi-Layer Perceptron (MLP) models, achieving an F-score of 0.73 and 0.69 for predicting valence and arousal values, respectively, on the MediaEval dataset. Sharma et al. [76] combined machine learning algorithms such as Support Vector Machines (SVM) and Naive Bayes (NB) to predict emotions using both audio and lyric features. They achieved an accuracy of 63% on the PMEmo dataset. Similarly, Griffiths et al. [92] developed a multi-genre MER model using linear regressors. They achieved a $R^{2}$ score of 0.776 and 0.85 for predicting valence and arousal values, respectively, on the self-built dataset. Meanwhile, Xia et al. [93] utilized clustering algorithms alongside machine learning techniques like SVM and K-Nearest Neighbors (KNN) for emotion recognition, and their hybrid model which combined all the machine learning techniques achieved an accuracy of 85% on the Solymani et al.’s dataset.
@@
A distinction between the various models can be made based on the input modality. Some models are exclusively based on MIDI such as the multi-task architecture proposed by [94], and as such use a token-based representation. Given the limited size of MIDI datasets, the accuracy of such models is limited, e.g., the model by [94] reaches 67.56% accuracy when predicting between four emotion classes. Most of the existing MER models are based on audio, and hence they take as input raw audio. This is often converted into Mel-spectrograms which are then processed through convolutional neural networks (CNNs) [83], or the audio could be directly processed through a WaveNet architecture, which is a type of temporal CNN [113]. In recent years, audio embeddings pretrained on large-scale datasets have become available, providing researchers with tools to enable transfer learning. For instance, [114] incorporated a VGGish network with Transformers, achieving mean squared error (MSE) and Pearson correlation coefficient (PCC) values of 0.117 and 0.655 for arousal, respectively, and 0.170 and 0.575 for valence, respectively.
Alternatives to these approaches include directly extracting spectral features (e.g., MFCCs, spectral centroids) with libraries such as OpenSmile, which has a configuration file specifically for the emotion recognition task [115]. [71] uses this approach and achieves $R^{2}$ scores of 0.378 and 0.638 for predicting dynamic valence and arousal values, respectively, on the MediaEval dataset.
@@
Recognizing the importance of induced emotions, recent research has focused on models that leverage physiological data to predict music-induced emotions. For instance, [44] constructed the HKU956 dataset with aligned peripheral physiological signals (i.e., heart rate, skin conductance, blood volume pulse, skin temperature) and self-reported emotion from 30 participants. The study revealed that physiological features significantly contribute to valence classification and that multimodal classifiers outperform single-modality ones. [89] uses PMEmo as induced emotion recognition dataset. This study merges audioLIME, a source-separation-based explainable model, with mid-level perceptual features to form an intuitive connection between input audio and emotion predictions, providing insights into model predictions. [103] also uses PMEmo as induced emotion recognition dataset and introduces a novel method named Modularized Composite Attention Network (MCAN). This method enhances feature extraction and employs attention mechanisms to improve the stability and accuracy of emotion prediction models.
@@
Identifying the current state-of-the-art Music Emotion Recognition (MER) model is challenging due to factors such as differences in datasets and performance metrics, as highlighted in the previous section. One standard evaluation is provided by the 2021 Emotion and Theme Recognition in Music competition, which uses the MTG-Jamendo dataset. The best performing model achieved a PR-AUC-macro of 0.150872 with a convolutional neural network (CNN) architecture [75].
Figure 1 shows which datasets have most been used in Music Emotion Recognition models since 2020, as presented in Table II. As shown, self-built datasets, which are created by the authors themselves for specific research purposes, are the most commonly used, appearing in 11 instances. These datasets may contain licenced data and are not available to the research community. Other frequently utilized datasets include DEAM (8 occurrences) and MTG-Jamendo (6 occurrences), indicating their popularity in MER research. Some datasets, such as PMEmo and Solymani et al.’s dataset, appear less frequently but still contribute valuable insights.
@@ V Challenges and future directions
Whereas the first publications on music and emotion surfaced in the 1930s [58], current MER models still struggle to match human performance in emotion recognition. The field of MER still faces several challenges and various opportunities for future exploration remain, ranging from overcoming data limitations to the integration of emerging technologies. Understanding and addressing these challenges is crucial for advancing the field and unlocking its full potential.
Dataset limitations One of the primary challenges in MER is the scarcity of large, diverse, copyright-cleared, emotion-annotated datasets. Limited datasets hinder the development and evaluation of robust MER models, leading to potential biases and generalization issues. We have recently seen advances in this area with the release of larger datasets such as MTG-Jamendo [36], Music4all [41], MuSe [43], which contain 18k, 109k, and 90k instances respectively. Whereas MTG-Jamendo and MuSe are available under a Creative Commons licence, however, Music4all contains copyrighted tracks. In addition, when exploring datasets, we notice that they are often skewed towards one particular genre. For instance, the DEAM dataset [32] consists mostly of rock and electronic music genres, while the VGMIDI dataset is focused solely on video game soundtracks.
@@
Finally, it is also hard to make comparisons across datasets, as many of them use different emotion representations. While some of these differences may be bridged by using synonyms, it is currently not a common practice to compare across datasets. However, if we are able to merge different emotion representations and train across multiple datasets, then we can start leveraging larger-scale models. Bridging between datasets that use continuous representations and categorical models is the least evident. [133] have identified a mapping between arousal/valence and categorical labels. This has been used by [6] to facilitate emotion-controlled lead sheet generation, and it may offer a way to merge continuous and categorical datasets into a larger-scale dataset for music emotion.

### v3

@@ IV Models and approaches
In this section, we briefly touch upon some of the more recent MER models, highlighting the state-of-the-art approaches over the last few years. This is not an exhaustive overview, and for a more comprehensive review, readers are referred to other survey papers [12, 13, 14, 10, 15]. The aim of this section is to point out the current state-of-the-art and various approaches over the last five years. The models were selected through Google Scholar searches using the search terms ‘music emotion prediction model’, ‘affective music prediction model’, and ‘music emotion recognition model’ starting from the year 2020, as well as by following links in the articles.
@@
Figure 2: Number of Music Emotion Recognition Models since 2020 that use the listed datasets.
@@
In recent years, various deep-learning MER models have been developed, employing techniques such as Convolutional Neural Networks (CNNs) [88, 90, 92, 93, 94, 83, 103, 96, 97, 98, 109, 116, 55, 112, 119, 122, 121], Recursive Neural Networks (RNNs) such as Gated Recurrent Units (GRUs) [112] and Long-Short Term Memory networks (LSTMs) [88, 89, 90, 91, 96, 97, 99, 109, 51, 55, 118], and more recently, Transformer architectures [100, 117, 55, 114].
Additionally, generative models such as Generative Adversarial Networks (GANs) have been applied. For instance, Huang et al. [102] developed a GAN-based model for emotion recognition using audio inputs under IoT environments.
@@
On the MTG-Jamendo dataset [41], several models have demonstrated noteworthy performance. For example, Mayerl et al. [84] employed a CNN-based (ResNet18) model using clustering-based ensembles, achieving a PR-AUC-macro score of 0.1087 and a ROC-AUC-macro score of 0.7047. Tan et al. [85], using a semi-supervised learning method and a CRNN architecture, reported higher scores of PR-AUC-macro 0.1357 and ROC-AUC-macro 0.7687. Pham et al. [98] achieved similar results with an EfficientNet-based CNN model, utilizing a co-teaching strategy to manage noisy data, resulting in PR-AUC-macro 0.1435 and ROC-AUC-macro 0.7599. Bour et al. [86] leveraged frequency-dependent convolutions in a CNN model, achieving PR-AUC-macro 0.1509 and ROC-AUC-macro 0.7748. Kang and Herremans [124] proposed a unified multitask learning framework that leverages both categorical and dimensional emotion labels by combining musical features (e.g., key and chords) with MERT representations. Through knowledge distillation across multiple datasets, this approach enhances cross-dataset generalization and achieves improved performance on MTG-Jamendo with a PR-AUC of 0.1543 and ROC-AUC of 0.7810. Knox et al. [95] proposed an ensemble of CNN-based models trained on Mel spectrograms, exploring the impact of different loss functions and resampling strategies for multi-label music tagging. They showed that using focal loss effectively addressed class imbalance, and adjusting the CNN’s receptive field further improved performance. Their model achieved the highest reported scores in the MediaEval “Emotion and Theme Recognition in Music” task, with a PR-AUC-macro of 0.1610 and ROC-AUC-macro of 0.7810.
@@
Notably, hybrid architectures combining multiple deep learning techniques, such as CNN-LSTM models, have shown promising results. For instance, Chen and Li [90] employed a CNN-LSTM model for emotion recognition using the Last.fm subset of the Million Song Dataset—a large-scale collection of music metadata with listener-generated tags. Their model outperformed standalone CNN and LSTM baselines, achieving an accuracy of 68.1% for audio classification and 74.2% for lyric classification. In contrast to classification approaches, He et al. [88] proposed a regression-based model using multi-view CNNs and bidirectional LSTMs directly on raw audio. Without relying on handcrafted features, their model achieved strong performance with RMSE scores of 0.219 for valence and 0.212 for arousal on the Solymani et al.’s dataset [28]. Transformer-based architectures have also gained traction in recent years, with notable examples by Suresh et al. [117] and Agrawal et al. [100], who explored their effectiveness in MER. Suresh et al., for example, applied a Transformer model to the MoodyLyrics dataset [35], which contains 2,595 tracks labeled with four mood categories derived from lyrics. Their model achieved the highest reported accuracy of 77.94%, outperforming CNN and Bi-GRU baselines.
@@
Other models have employed more traditional machine learning techniques alongside deep learning approaches. Medina et al. [110] focused on classifying emotions from audio using SVM, Random Forest, and Multi-Layer Perceptron (MLP) models, achieving an F-score of 0.73 and 0.69 for predicting valence and arousal values, respectively, on the DEAM dataset [37]. Sharma et al. [87] combined machine learning algorithms such as Support Vector Machines (SVM) and Naive Bayes (NB) to predict emotions using both audio and lyric features. They achieved an accuracy of 63% on the PMEmo dataset [38]. Similarly, Griffiths et al. [104] developed a multi-genre MER model using linear regressors. They achieved a $R^{2}$ score of 0.776 and 0.85 for predicting valence and arousal values, respectively, on the self-built dataset. Meanwhile, Xia et al. [105] utilized clustering algorithms alongside machine learning techniques like SVM and K-Nearest Neighbors (KNN) for emotion recognition, and their hybrid model which combined all the machine learning techniques achieved an accuracy of 85% on the Solymani et al.’s dataset [28].
@@
A distinction between the various models can be made based on the input modality. Some models are exclusively based on MIDI such as the multi-task architecture proposed by [106], and as such use a token-based representation. Given the limited size of MIDI datasets, the accuracy of such models is limited, e.g., the model by [106] reaches 67.56% accuracy when predicting between four emotion classes. Most of the existing MER models are based on audio, and hence they take as input raw audio. This is often converted into Mel-spectrograms which are then processed through convolutional neural networks (CNNs) [94], or the audio could be directly processed through a WaveNet architecture, which is a type of temporal CNN [127].
In recent years, audio embeddings pretrained on large-scale datasets have become increasingly common, enabling transfer learning for Music Emotion Recognition (MER). Beyond traditional embeddings, recent research has explored both supervised and unsupervised pretraining strategies for audio representation learning, where MER is treated as a downstream task. For instance, MERT (Music undERstanding model with self-supervised Training) [128] is specifically tailored to music, addressing challenges like pitch and tonality through a novel training scheme. MERT incorporates pseudo labels from two teacher models—an acoustic teacher based on Residual Vector Quantization Variational AutoEncoder (RVQ-VAE) and a musical teacher based on the Constant-Q Transform (CQT)—within a masked language modeling framework. When used as a feature extractor for MER, MERT-95M achieves a PR-AUC of 0.1340 and ROC-AUC of 0.7640 on the MTG-Jamendo dataset [41], while MERT-330M reaches a PR-AUC of 0.1400 and ROC-AUC of 0.7650. Building upon MERT embeddings, Kang and Herremans [124] introduce a unified multitask learning framework that leverages both categorical and dimensional emotion labels. Their model achieved a PR-AUC of 0.1543 and ROC-AUC of 0.7810 on the MTG-Jamendo dataset [41]. For dimensional emotion regression, it attained $R^{2}$ scores of 0.5473 for valence and 0.7940 for arousal on PMEmo [38], 0.5184 for valence and 0.6228 for arousal on DEAM [37], and 0.6512 for valence and 0.7616 for arousal on the dataset by Solymani et al. [28].
McCallum et al. [129] present a comparative analysis of audio representation learning strategies, showing that supervised pretraining on expert-annotated music datasets leads to state-of-the-art performance in emotion tagging. Their models achieve 78.6 ROC-AUC and 16.1 PR-AUC on the MTG-Jamendo Mood/Theme dataset—outperforming previously reported benchmarks. Alonso-Jiménez et al. [130] propose MAEST, a transformer-based architecture trained with patchout and pre-initialized with ImageNet or AudioSet weights. Although its performance is slightly lower (78.1 ROC-AUC and 15.4 PR-AUC), MAEST demonstrates the potential of efficient, convolution-free architectures in music emotion tagging, particularly when speed and scalability are important.
@@
Alternatives to these approaches include directly extracting spectral features (e.g., MFCCs, spectral centroids) with libraries such as OpenSmile, which has a configuration file specifically for the emotion recognition task [131]. [82] uses this approach and achieves $R^{2}$ scores of 0.378 and 0.638 for predicting dynamic valence and arousal values, respectively, on the Solymani et al.’s dataset [28]. Recognizing the importance of induced emotions, recent research has focused on models that leverage physiological data to predict music-induced emotions. For instance, [49] constructed the HKU956 dataset [49] with aligned peripheral physiological signals (i.e., heart rate, skin conductance, blood volume pulse, skin temperature) and self-reported emotion from 30 participants. The study revealed that physiological features significantly contribute to valence classification and that multimodal classifiers outperform single-modality ones. [101] uses PMEmo [38] as induced emotion recognition dataset. This study merges audioLIME, a source-separation-based explainable model, with mid-level perceptual features to form an intuitive connection between input audio and emotion predictions, providing insights into model predictions. [115] also uses PMEmo as induced emotion recognition dataset and introduces a novel method named Modularized Composite Attention Network (MCAN). This method enhances feature extraction and employs attention mechanisms to improve the stability and accuracy of emotion prediction models.
@@
Identifying the current state-of-the-art Music Emotion Recognition (MER) model is challenging due to factors such as differences in datasets and performance metrics, as highlighted in the previous section. A commonly used benchmark is the Emotion and Theme Recognition in Music competition based on the MTG-Jamendo dataset, where the top-performing model by Knox et al. [95] achieved a PR-AUC-macro of 0.161 and a ROC-AUC-macro of 0.781 using an ensemble of CNN-based models trained with focal loss and receptive field tuning. While benchmark performances of MER models have steadily improved (e.g., some achieving ROC-AUC scores over 0.78 on datasets like MTG-Jamendo), their translation to real-world applications remains limited.
One area where we do see emotion models integrated is emotion-conditioned music generation, where preliminary MER models are incorporated to guide the emotional content of generated music (e.g., Makris et al. [6]). Furthermore, emotion detection systems have been piloted in adaptive gaming environments or personalized music therapy applications [5], but large-scale real-world deployment of deep MER models remains rare due to variability across listeners, subjective annotation challenges, and differences between training and inference conditions.
Figure 2 illustrates the most frequently used datasets by the listed Music Emotion Recognition (MER) models since 2020, as summarized in Table II. This count includes studies where MER is the primary task as well as those where it serves as a downstream application, such as in music representation learning [128]. Self-built datasets, created by authors for specific research purposes, remain the most commonly used, appearing in 11 instances. These datasets often contain licensed music, restricting their accessibility to the broader research community. Other frequently utilized datasets include DEAM (8 occurrences) and MTG-Jamendo (7 occurrences), highlighting their importance in MER studies. Some datasets, such as PMEmo [38] and Solymani et al.’s dataset [28], appear less frequently but still contribute valuable insights. For instance, PMEmo [38] includes dynamic emotion labels and physiological signals (EDA), enabling multimodal affective analysis, while Solymani et al.’s dataset [28] offers continuous valence-arousal annotations and standard deviations, supporting studies on annotation reliability and temporal emotion dynamics.
@@ V Challenges and future directions
Whereas the first publications on music and emotion surfaced in the 1930s [65], current MER models still struggle to match human performance in emotion recognition. The field of MER still faces several challenges and various opportunities for future exploration remain, ranging from overcoming data limitations to the integration of emerging technologies. Understanding and addressing these challenges is crucial for advancing the field.
Dataset limitations One of the primary challenges in MER is the scarcity of large, diverse, copyright-cleared, emotion-annotated datasets. Limited datasets hinder the development and evaluation of robust MER models, leading to potential biases and generalization issues, and also prevent the establishment of reliable performance benchmarks across studies.
We have recently seen advances in this area with the release of larger datasets such as MTG-Jamendo [41], Music4all [46], MuSe [48], which contain 18k, 109k, and 90k instances respectively. Whereas MTG-Jamendo and MuSe are available under a Creative Commons licence, however, Music4all contains copyrighted tracks. In addition, when exploring datasets, we notice that they are often skewed towards one particular genre. For instance, the DEAM dataset [37] consists mostly of rock and electronic music genres, while the VGMIDI dataset is focused solely on video game soundtracks.
@@
While there have been some efforts toward cross-dataset comparisons, such practices are still relatively rare. A notable example is the MusAV dataset [53], which provides a benchmark for evaluating arousal-valence (AV) regression models trained on different datasets. MusAV uses relative pairwise comparisons as ground truth and enables comparative validation across models trained on diverse AV datasets. Such initiatives are promising steps toward better generalization and standardized evaluation, but they are currently exceptions rather than the rule. In general, many datasets still use different emotion representations, and cross-dataset evaluation remains challenging. Recent efforts have aimed to address this limitation: Kang et al. [124] propose a unified multitask framework to combine categorical and dimensional labels, while Liu et al. [123] use LLM-based label embeddings to align emotion annotations across datasets and enable zero-shot generalization.
Bridging between datasets that use continuous representations and categorical models is not straightforward. Several studies, such as Paltoglou and Thelwall [149], have proposed mappings between arousal/valence and categorical labels. This approach has been further utilized, for example, by Makris et al. [6] for emotion-controlled lead sheet generation. In addition, large-scale affective norm resources such as Warriner et al. [150], which provides valence, arousal, and dominance (VAD) ratings for 13,915 English lemmas, and Mohammad [151], who developed the NRC VAD Lexicon with ratings for over 20,000 words, offer valuable foundations for linking continuous and categorical representations. Such mappings could support merging continuous and categorical datasets into larger-scale resources for music emotion research.
@@
Toward reliable and comparable MER research Despite significant advancements, the field still lacks standardized, high-quality datasets that serve as universal benchmarks. As discussed above, most existing datasets differ widely in genre coverage, emotion representation models, annotation procedures, and data quality. This heterogeneity, combined with inconsistent use of evaluation metrics and train/test splits, often makes it difficult to compare results across studies—even when using the same dataset. For example, the MTG-Jamendo dataset has been used in numerous studies (see Table II), yet reported results vary substantially due to differing preprocessing strategies, loss functions, or model inputs. This wide variance raises questions about reproducibility and comparability in the field.
While recent initiatives like MusAV [53] and unified multitask frameworks [124] represent important progress, the lack of standardized benchmarks still hinders progress. We argue that no single dataset can serve all research needs, especially given the subjective and culture-specific nature of emotion. Rather than searching for a one-size-fits-all dataset, we encourage the development of dataset-agnostic benchmarking tools, clearly defined splits, and model evaluation protocols. The community would benefit from centralized efforts (e.g., leaderboards, open splits, and documentation hubs) that promote transparency and make it easier to assess which models and results should be taken more seriously. Without such measures, the field risks being undermined by non-comparable results and hard-to-replicate studies.